what is the area and perimeter of the largest stained glass window in the world, if it's 8m high and 95 m long?
Answer by jeseca_1964@yahoo.com(2)

A = L W
= 8 (95)
= 760 sq.m
P = 2(L)+ 2(W)
= 2(95) + 2(8)
= 190 + 16
= 216m